demonstrate how to write an equation(3,-2) and perpendicular to the line y=3x-4
Answer by ewatrrr(24785)


Hi
line that passes through (3, -2) and is perpendicular to y=3x-4

y=3x-4 |slope m = 3
perpendicular lines have slopes that are neagtive reciprocals of one another
New Line: m = -1/3 |Note: 3*-1/3 = -1
y = (-1/3)x + b |using ordered pair Pt(3,-2) to solve for b
-2 = -1 + b
-1 = b
y = (-1/3)x - 1 |written in slope-intercept form